OpenStudy (anonymous):

Glenn needs 20 minutes to clean his room. If his little sister Veronika pulls things out while he is cleaning, it takes 60 minutes for him to clean the room. How many minutes does it take for Veronika to pull everything out? 10 minutes 12 minutes 30 minutes 50 minutes

OpenStudy (anonymous):

im going with b

OpenStudy (anonymous):

i have seen this before, and actually figured out what you are supposed to say i am sure we can do it

OpenStudy (anonymous):

yeah i remember now

OpenStudy (anonymous):

solve \[\frac{1}{20}-\frac{1}{x}=\frac{1}{60}\]

OpenStudy (anonymous):

i would say " takes her two seconds to pull everything out" because this is one of the sillier questions i have ever see, but if you solve the above equation you get \(x=30\)

OpenStudy (anonymous):

i know i am not supposed to give a direct answer but this question is so foolish that there is no point in an explanation